# machine learning, optical communications, passive optical network, equalization, optical performance monitoring, modulation format identification, bit error rate, optical signal-to-noise ratio, nonlinearities

### v1

@@ 3 DSP for Signal Equalization in Communication Systems
In this section, we provide an overview of conventional signal equalization techniques, ranging from basic zero-forcing equalization to more advanced approaches such as feed-forward equalizers (FFE), decision-feedback equalizers (DFE), Viterbi and Volterra equalizers, as well as adaptive equalizers. We discuss the advantages and limitations of these techniques and present a comparative analysis alongside ML models. We briefly introduce the principles of these methods and provide a complexity analysis for each of them in combination with traditional sequential ML methods in the Table 1.
@@
Zero Forcing (ZF): Zero forcing is an LE derived by minimizing inter-symbol interference (ISI). A groundbreaking study Tufts (1965) has established the analytical foundation for optimal zero-forcing and minimum mean squared error (MSE) equalization in channels with additive white noise and specified frequency response. The study further demonstrates that an optimal LE can be implemented as a cascade of filters, with taps spaced at symbol intervals. However, when the channel effect exhibits a deep frequency response “valleys,” equalization will yield poor performance due to the noise enhancement problem.
@@
Feed Forward Equalizer: FFE RagaLasya (2017) mitigates ISI in communication channels by processing the received signal forwardly without feedback. Its simplicity makes it suitable for systems where feedback is challenging or undesired.
@@
Decision Feedback Equalizers: Due to the noise enhancement, DFE is designed to reduce ISI by subtracting already-known symbols. In this way, ISI from already detected symbols is eliminated. Adaptation of the forward and feedback filters of DFEs follow the same pattern as for LEs WilliamsonD (1992). The disadvantage is that it could induce accumulated errors associated with feeding back filtered incorrect detector decisions.
@@
Viterbi equalizer: The Viterbi equalizer seeks to estimate the most likely sequence of transmitted symbols, given the received sequence. By constructing a trellis diagram where nodes represent possible transmitted symbols and transitions denote potential channel transitions, the Viterbi algorithm dynamically optimizes path metrics to identify the most probable sequence. This process involves state transition probabilities and meticulous calculations to mitigate the impact of channel impairments. Mathematically, the Viterbi equalizer applies the Viterbi algorithm, which belongs to the dynamic programming algorithm for finding the most likely sequence of hidden states in a hidden Markov model. The time complexity of the Viterbi equalizer is determined by the Viterbi algorithm, which is influenced by the length of the input sequence and the number of states, making it $O(T\cdot N^{2})$, where $T$ denotes the length the sequence and $S$ refers to the number of the hidden states Forney (2005).
@@
Volterra equalizer: Is a nonlinear equalizer used in optical communication systems to compensate for nonlinear distortions introduced by the fiber channel volterra (2103). In the context of PAM4 optical systems, a Volterra equalizer can be employed to mitigate the nonlinear effects that arise due to the intensity-dependent response of the optical components. The Volterra equalizer is based on the Volterra series expansion, which allows for the modeling of nonlinear systems. The key idea is to use a set of nonlinear filters, known as Volterra kernels, to capture the nonlinear characteristics of the channel. In the case of PAM4 optical systems, the Volterra equalizer is typically designed to handle the specific nonlinearities associated with the modulation format and the optical components employed. The structure of a Volterra equalizer consists of multiple stages, each representing a different order of nonlinearity. The first stage corresponds to the linear equalizer, which performs initial equalization to address linear distortions. Subsequent stages of the Volterra equalizer capture and compensate for higher-order nonlinear distortions. These stages involve nonlinear filters that take multiple past symbols as inputs and produce outputs based on their interaction. The number of stages and the complexity of the Volterra equalizer depend on the specific system requirements and the level of nonlinear distortions present. The coefficients of the Volterra kernels are typically adapted or optimized using algorithms such as the least mean squares (LMS) or recursive least squares (RLS) algorithms. These algorithms iteratively adjust the coefficients based on the error between the equalized signal and the desired signal, aiming to minimize the distortion and improve the overall system performance.
@@ 4 Traditional Sequential ML Methods
With the increasing demand for higher data transmission rates and the limitations of traditional prediction methods reaching their practical limits in terms of accuracy, the need for algorithms with high precision, reliability, and low complexity has become urgent. In this section, we introduce a solution in DL to address this challenge. In summary, we begins with an overview of relevant research studies, subsequently providing a chronological exploration of key sequential models, namely, recurrent neural network (RNN), long short-term memory (LSTM), gated recurrent unit (GRU), and convolutional neural network (CNN). These architectures, which serve as fundamental components in DL models, are systematically introduced, along with concise examples that clarify their design principles, application methodologies, and complexity analysis.
@@
In 2018, Karanov et al. BKaranov (2018) introduced an end-to-end deep neural network system for optical communications, encompassing the entire chain of a transmitter, receiver, and channel model. This pioneering work demonstrated the feasibility of transceiver optimization in an end-to-end manner. Owing to the intrinsic sequential structure of communication systems, sequential models, including LSTM networks GravesA (2012), RNNs PascanuR (2013), and GRUs JingL (2017) have been extensively employed since their inception. They continue to serve as crucial baseline models and foundational elements for developing more efficient algorithms.
@@
RNN: It is originally designed for machine translation in natural language processing. This model is based on the Markov assumption about the hidden state and output sequence: namely, the output sequence depends only on the current potential state $h_{t}$. The potential state depends on the previous moment’s latent $h_{t-1}$ and input variables $x_{t-1}$ rather than on the historical data $\{x_{(t-1,\dots,0)},h_{(t-1,\dots,0)}\}$. Renowned for their adaptability in handling variable-length sequences and preserving state information across elements, RNNs find valuable applications in diverse communication fields XieY (2022). In recent work, RNN has shown promising results in equalization than the previous joint methods based on Volterra and Viterbi equalizers in two-dimensional eight-level pulse amplitude modulation (2D-PAM8) links X. Qin (2023). Despite the promising performance, it suffers from exploding gradient issues caused by the direct gradient flow of multiple layers Pascanu (2012). In other words, when the largest singular value of the gradient matrix with respect to parameters is larger than $1$ in the backpropagation process, the numerical instability will result in the training difficulty IEEE (1994).
LSTM: In contrast, LSTM addresses this issue by introducing input and forget gates to determine the gradient flow jointly. These additional states facilitate the learning of a weight matrix that preserves useful information in the hidden state Pascanu (2012). Here, we briefly recapitulate the architecture of LSTM; for more details, please refer to the original publication Sepp Hochreiter (1997). The input gate decides the information from the current input stored in the remote state, the forget gate regulates the retention or discarding of data from the prior hidden state, and the output gate defines the information presented as the final prediction. This gating mechanism equips LSTMs with the ability to effectively preserve pertinent details from lengthy sequences while discarding extraneous information. It is noteworthy, however, that LSTM does not guarantee a solution without vanishing/exploding gradients. Still, it does provide a more accessible way for the model to learn long-distance dependencies. LSTM has been widely applied in short-reach communication, especially for tasks that require the ability to capture intricate long-term dependencies LingP (2022).
@@
GRU: Similar to LSTM, GRU is widely utilized as a variant of RNNs. It simplifies four hidden states into two critical gates, namely the update gate and the reset gate DeyR (2017). The update gate determines the relevant information to retain from the prior hidden state and the current input, while the reset gate dictates the data to discard. The resultant hidden state synthesizes the information preserved by the update gate and the current input. From an application perspective, GRUs find utility in scenarios where temporal dependencies and adjacent symbols within sequences play a crucial role. For instance, in short-reach communication, GRUs can alleviate signal distortions caused by chromatic dispersion and nonlinearities StavrosD (2021). For example, recent work has shown that the quality(Q)-factors can exceed the hard-decision forward error correction (HD-FEC) limit of 8.52 dB with the aid of bi-GRU nonlinear equalizer (8.52 dB estimated from $Q=20\log_{10}(\sqrt{2}\text{erfc}^{-1}(2\text{BER}))$)  LiuX (2021) .
CNN: Although CNN does not strictly belong to sequence models, it is widely applied in various domains due to its high parameter efficiency, weight-sharing mechanism, and plug-and-play characteristics NWu (2019). Convolutional operations use a convolutional kernel to scan the input signal with a fixed step on a specific dimension to capture the temporal features that are important for the downstream task. Typically, it is succeeded by a pooling layer and a nonlinear activation function, serving as redundancy reduction and nonlinear transformation, respectively. Essentially, features extracted by the convolutional kernel that closely align with the input signal are more likely to be preserved in the weight matrix. The relevant features crucial for downstream tasks will be enhanced by backpropagation processing. Recently, CNN has been utilized alongside belief propagation (BP) to improve decoding performance in an iterative manner Zhang (2023). Furthermore, a large number of test cases have shown that the combination of multiple layers of small convolutional kernels is often more efficient than using large kernels Szegedy (2014). Widely used and accepted foundational modules include inception and inception reduction, which extract temporal dependencies of different scales by employing concatenation of $1\times 1$, $3\times 3$, and $5\times 5$ convolutional kernels. In addition, a convolution kernel in the shape of $1\times 1$ serves a special function, thus reducing the channel dimension. It is generally used between two convolution layers or the output layer.
Summary: In this section, we have provided an introduction to the most common units of ML in optical communication systems. As of today, they are still directly used. In light of this, we summarize the aforementioned models’ complexity as follows. We neglect the choice of epochs and batch size but provide the complexity analysis for each batched sample.
@@ 5 Advanced Sequential ML Methods
In Section 4, we introduced traditional sequence models, such as RNN, LSTM, GRU, and CNN. The key question we aim to answer in this section is how to effectively incorporate the unique characteristics of time series data into the modeling process and leverage the temporal convolution model to mitigate channel distortion. Different from other deep learning models to be introduced in the following sections, such as a transformer and Fourier-based Neural Networks, which necessitate meticulous parameter initialization and hyperparameter tuning when applied to new datasets, convolutional models exhibit superior generalization performance and are robust to parameter perturbation LIU (2022).
This section starts with channel modeling, encompassing four distinct noise models. We derive the characteristics and capabilities required for the algorithm based on these models. Subsequently, we provide a detailed exposition of the architectures and fundamental assumptions underlying three models: Frequency-Calibrated Sampled-Interaction Neural Network (FC-SCINet) CShao (2014)), DLinear Zeng (2022) and LightTS ZhangT (2022).
5.1 Channel Modelling:
The main limiting factor for the equalization task in a short-reach/PON system is ISI as a result of Chromatic Dispersion (CD), sampling error (Jitter), frequency shift (chirp), and other nonlinear effects BKaranov (2018).
Chromatic Dispersion in an optical communication system is caused by different phase velocities with respect to frequency. It fundamentally constitutes a linear transformation, and its mathematical representation involves a differential equation that considers spatial position and time, which can be presented as
@@ 5.1 Channel Modelling:
Jitter typically manifests as high-frequency fluctuations in the amplitude of an individual signal, and this elevated-frequency perturbation constitutes a highly effective form of attack on neural networks dependent on low-frequency signals.
@@
Figure 1: Sampling modules in temporal convolution networks modified from CShao (2014); ZhangT (2022)
@@ 5.2 Temporal Convolution Neural Network
Deep learning-based equalizers fundamentally capture intricate, domain-specific nonlinear disturbances by employing linear transformations and nonlinear activation functions. Within the temporal convolutional neural network, the core modules comprise interval, continuous, and interaction sampling modules, alongside convolutional neurons and linear layers, as depicted in Fig. 1. Each of these modules offers practical flexibility for hardware implementation, due to their computational efficiency. Subsequently, we delve into three of the most efficient convolution-based sampling networks.
@@
SCIBlock: The SCIBlock, a crucial component of FC-SCINet, exhibits remarkable capabilities as it iteratively decomposes a signal into sub-sequences at various scales while incorporating nonlinear transformations between adjacent layers. In contrast, the decomp block is restricted to enhancing fixed signal components and is limited to a single scale. From a mathematical perspective, the SCIBlock applies a hierarchical structure by systematically downsampling the input sequence into even-positioned and odd-positioned samples, denoted as $x_{even}$ and $x_{odd}$. Following the convolutional layer, the sub-sequences in adjacent layers are iteratively multiplied together utilizing exponential and multiplication operations, as shown in Eq. (10) and (11).
@@
DLinear: As previously mentioned in FC-SCINet, while the concatenation of the decomp module may not offer optimal equalization, it has exhibited superior performance in real-world datasets. Therefore, we will provide a brief introduction to this module. It first decomposes a raw data input into a low $\mathbf{x_{s}}$ and high frequency $\mathbf{x_{f}}$ signal. $\mathbf{x_{s}}$ is extracted by a moving average kernel. It is equivalent to filtering the signal using a $sinc$ function in the frequency domain. These two components are added in a linear combination form, expressed by $W_{s}$, $W_{f}$. The operation above is presented in Eq.(12).
@@
By iterative decomposition with different kernel sizes, DLinear can be extended to a deeper network. The complexity is $O(kn_{s})$, where $k$ is the number of the layer, $n_{s}$ is the length of the model input. To simplify the complexity, the weight matrix $W$ could be replaced by the convolutional kernel.
@@
Sampling: In contrast to SCIBlock, which samples the raw sequence using odd and even indices, LightTS introduces two generic sampling strategies: Interval Sampling and Continuous Sampling, as shown in Fig. 1. Interval sampling partitions time series data into non-overlapping sub-sequences based on fixed time intervals, as shown in Fig 1. This approach helps identify periodic patterns or regularities within the data while minimizing information loss. On the other hand, continuous sampling divides sequences into continuous sub-sequences, extracting data points continuously throughout the time series and preserving temporal continuity. This sampling method enables the capture of patterns within the period, ensuring a more comprehensive representation of the underlying dynamics. The subsequent section presents an MLP-based architecture to extract useful features from both the down-sampled sub-sequences and continuously sampled sub-sequences.
@@
IEBlock: The IEBlock serves as the central module in LightTS, designed to effectively process the 2D matrix resulting from continuous sampling and interval sampling. This block comprises three essential components: 1) temporal projection, which identifies temporal features following continuous sampling; 2) channel projection, which captures inter-channel information following interval sampling; and 3) the exchange block, which integrates the information from the outputs mentioned above, facilitating information fusion, all of them utilize MLP as the nonlinear behavior learning module. The design of LightTS is notably concise, employing only two sampling modules and an MLP. On certain datasets, it surpasses the performance of FC-SCINet Zeng (2022).
@@
The above models require structural adaption, additional preprocessing etc. while applied in the practise. One of the recent work which has been successful in PONs applications includes the FC-SCINet CShao (2014). Different from LSTM, which offers the advantage of ensuring information flows strictly from past to future, temporal CNN goes beyond this by modeling the global dependency between input and output, while also leveraging stacked causal convolution layers. Additionally, the FC-SCINet introduces interaction modeling, enabling the explicit capture of interactions between elements within a sequence, making it a more advanced alternative to LSTM. In addition to these benefits, convolutional CNNs and SCInet offer several advantages over LSTM:
@@
  • Convolutional CNNs can identify patterns regardless of their position within the input sequence. This property makes them well-suited for tasks where the position or timing of relevant features is not fixed, providing greater flexibility compared to LSTM.
@@
  • Convolutional CNNs excel at capturing local patterns and extracting relevant features from the input sequence. This ability is particularly useful for tasks that require identifying and recognizing specific patterns or motifs within the data.
@@
  • Convolutional CNNs and SCInet architectures typically have fewer parameters compared to LSTM models. This reduced parameter count can make training and inference more efficient, especially when working with limited computational resources or when dealing with large datasets.
@@
Recent Progress: Convolutional networks maintain an indispensable role in current time series prediction research and practical applications due to their high parameter efficiency, model stability, and robust theoretical foundation (Multiscale Decomposition). The complexity of these networks primarily relies on the number of layers and the size of the convolutional kernels. Presently, advanced designs like dilated convolution and inception are frequently integrated with other modules to construct sophisticated deep learning models, yet they are seldom employed independently. Nevertheless, temporal convolutional neural networks retain an indispensable edge in terms of the Performance-vs-Complexity ratio and prove straightforward to implement in hardware.

### v2

@@ 3 DSP for Signal Equalization in Communication Systems
In this section, we provide an overview of conventional signal equalization techniques, ranging from basic zero-forcing equalization to more advanced approaches such as feed-forward equalizers (FFE), decision-feedback equalizers (DFE), Viterbi and Volterra equalizers, as well as adaptive equalizers. We discuss the advantages and limitations of these techniques, comparing the performance of ML models. Table 1 provides the complexity analysis for each method.
@@
Zero Forcing: It is a linear equalizer (LE) derived by minimizing inter-symbol interference (ISI). A study in Tufts (1965) has established the analytical foundation for optimal zero-forcing and minimum mean squared error (MSE) equalization in channels with additive white noise and specified frequency response. The study demonstrates that an optimal LE can be implemented as a cascade of filters, with taps spaced at symbol intervals. However, when the channel effect exhibits a deep frequency response “valleys,” equalization will yield poor performance due to noise enhancement.
@@
Feed Forward Equalizer: FFE RagaLasya (2017) mitigates ISI in communication channels by processing the received signal forwardly without feedback. Its simplicity makes it suitable for systems where feedback is unstable or challenging for implementation.
@@
Decision Feedback Equalizers: Due to the noise enhancement, DFE is designed to reduce ISI by subtracting already-known symbols. In this way, ISI from already detected symbols is eliminated. Adaptation of the forward and feedback filters of DFEs follow the same pattern as for LEs WilliamsonD (1992). The disadvantage is that it could potentially lead to accumulated errors from feeding back incorrect detection decisions
@@
Viterbi equalizer: The Viterbi equalizer seeks to estimate the most likely sequence of transmitted symbols, given the received sequence. By constructing a trellis diagram where nodes represent possible transmitted symbols and transitions denote potential channel transitions, the Viterbi algorithm dynamically optimizes path metrics to identify the most probable sequence. This process involves state transition probabilities and precise calculations to mitigate the impact of channel impairments. Mathematically, the Viterbi equalizer applies the Viterbi algorithm, which belongs to the dynamic programming algorithm for finding the most likely sequence of hidden states in a hidden Markov model. The time complexity of the Viterbi equalizer is determined by the Viterbi algorithm, which depends on the length of the input sequence and the number of states, making it $O(T\cdot N^{2})$, where $T$ denotes the length of the sequence and $N$ refers to the number of hidden states Forney (2005).
@@
Volterra equalizer: This is a nonlinear equalizer used in optical communication systems to compensate for nonlinear distortions introduced by the fiber channel volterra (2103). In PAM4 systems, severe ISI can be introduced due to the imperfect bandwidth of optical and electrical components. The main bandwidth bottleneck in IMDD systems comes from the transmitter side, as the achievable bandwidth of receiver-side devices is typically twice as high as the bandwidth of transmitter-side devices. In such scenarios, the Volterra equalizer can be effectively employed to address both a potential nonlinearity from the transmitter and the bandwidth limitations of the optical components. The higher-order Volterra kernels can model the frequency-dependent distortion and nonlinear effects caused by the limited transmitter bandwidth and nonlinear devices, such as Mach-Zehnder modulators. The Volterra equalizer is based on the Volterra series expansion, which allows for the modeling of nonlinear systems. The key idea is to use a set of nonlinear filters, known as Volterra kernels, to capture the nonlinear characteristics of the channel. The structure of a Volterra equalizer consists of multiple stages, each representing a different order of nonlinearity. The first stage corresponds to the linear equalizer, which performs initial equalization to address linear distortions. Subsequent stages of the Volterra equalizer capture and compensate for higher-order nonlinear distortions. These stages involve nonlinear filters that take multiple past symbols as inputs and produce outputs based on their interaction. The number of stages and the complexity of the Volterra equalizer depend on the specific system requirements and the level of nonlinear distortions present. The coefficients of the Volterra kernels are typically adapted or optimized using algorithms such as the least mean squares (LMS) or recursive least squares (RLS) algorithms. These algorithms iteratively adjust the coefficients based on the error between the equalized signal and the desired signal, aiming to minimize the distortion and improve the overall system performance.
@@ 4 Traditional Sequential ML Methods
With the increasing demand for higher data transmission rates and the limitations of traditional prediction methods reaching their practical limits in terms of accuracy, the need for algorithms with high precision, reliability, and low complexity has become urgent. In this section, we introduce new DL-based models to address this challenge. We overview relevant research studies, providing a chronological exploration of key sequential models, namely, recurrent neural networks (RNN), long short-term memory (LSTM), gated recurrent unit (GRU), and convolutional neural networks (CNN). The key architectural parts of DL models are explained, with clear examples showing how they work, how they are used, and how complex they are.
@@
In 2018, Karanov et al. BKaranov (2018) introduced an end-to-end deep neural network system for optical communications, encompassing the entire chain of a transmitter, receiver, and channel model. This research showed that transceiver optimization can be done in a complete, end-to-end way. Owing to the sequential structure of communication systems, sequential models, including LSTM networks GravesA (2012), RNNs PascanuR (2013), and GRUs JingL (2017) have been extensively employed. They are considered as baseline algorithms in order to generate more advanced and efficient algorithms.
@@
RNN: Originally designed for machine translation in natural language processing, this model is based on the Markov assumption about the hidden state and output sequence: the output sequence depends only on the current potential state $h_{t}$. The potential state depends on the previous moment’s latent $h_{t-1}$ and input variables $x_{t-1}$ rather than on the historical data ${x_{(t-1,\dots,0)},h_{(t-1,\dots,0)}}$. Renowned for their adaptability in handling variable-length sequences and preserving state information across elements, these models find valuable applications in diverse communication fields XieY (2022). In recent work, they have shown promising results in equalization compared to benchmark methods based on Volterra and Viterbi equalizers in two-dimensional eight-level PAM (2D-PAM8) links X. Qin (2023).
Despite its great equalization performance, this model suffers from exploding gradient issues caused by the direct gradient flow of multiple layers Pascanu (2012). In such networks, the backpropagation of the gradient is done by accumulating the gradient matrix. This can cause the gradient to grow exponentially if the eigenvalues of the gradient matrix are greater than 1, making the training process very difficult to converge. Conversely, when the eigenvalues of the gradient matrix are less than 1, the gradient will decrease over time until it vanishes completely, causing the parameters to stop updating IEEE (1994).
LSTM: The LSTM architecture can assist in overcoming this issue by extending the hidden state to a cell state, which is built using a gating mechanism. This mechanism has input, forget, and output gates that help control the flow of information Sepp Hochreiter (1997). LSTM models have additional internal states beyond just the hidden state. This allows them to learn a weight matrix that can better preserve useful information in the hidden state. The input gate decides what new information from the current input to be stored in the cell state. The forget gate decides what memories from the previous cell state to keep or discard. The output gate controls what information gets passed to the next cell state. This gating mechanism provides the ability to effectively hold onto relevant details from long sequences, while filtering out irrelevant information. This makes it easier to learn dependencies between distant parts of the input. As a result, LSTMs have been widely used in short-range communication tasks that require capturing complex long-term relationships in the data LingP (2022).
@@
GRU: This architecture simplifies the gating mechanism used in LSTM models. It has an update gate and a reset gate, instead of the three gates in LSTM DeyR (2017). The update gate determines what relevant information to retain from the previous state and the current input. The reset gate controls what data to discard. It is useful in scenarios where the temporal dependencies and relationships between adjacent symbols in a sequence are important. For example, in short-range communication systems, GRUs can help mitigate signal distortions caused by CD and nonlinearities StavrosD (2021). Recent research in 120 Gb/s coherent 64-quadrature amplitude modulated optical systems for transmission at 375 km, has shown that using a bi-directional GRU as a nonlinear equalizer can help improve the quality factor (Q-factor) beyond the 8.52 dB limit (8.52 dB estimated from $Q=20\log_{10}(\sqrt{2}\text{erfc}^{-1}(2\text{BER}))$)  LiuX (2021), typically required for hard-decision forward error correction (HD-FEC).
CNN: CNNs are not technically considered sequential models. However, they are widely used across many different domains. This is because of its important advantages, such as high parameter efficiency, weight-sharing mechanism, and plug-and-play characteristics NWu (2019). CNNs use a convolutional kernel to scan the input signal in a specific dimension, capturing temporal features that are important for the task at hand. This convolutional layer is typically followed by a pooling layer and a nonlinear activation function. The pooling layer reduces redundancy, while the activation function introduces nonlinearity. The convolutional kernel is designed to extract features that closely match the input signal, afterwards, backpropagation is used to optimize the weights of the network. This allows the CNN to learn and enhance the features that are most relevant for the target task. The weights in the network’s weight matrix are updated through backpropagation to amplify the important features needed for effective performance on the given ML problem. Furthermore, it has been observed that using multiple layers of small convolutional kernels is often more efficient than using large kernels. This approach, known as the inception architecture, was first introduced in the GoogleNet model Szegedy (2014). Two commonly used blocks in CNN are the inception module and the inception reduction module, which extract temporal dependencies of different scales by employing a concatenation of $1\times 1$, $3\times 3$, and $5\times 5$ convolutional kernels. In addition, It also uses a special type of convolutional kernel with a size of $1\times 1$. This $1\times 1$ convolution serves a unique purpose - it helps to reduce the number of feature map channels or dimensions. It is commonly used between two regular convolution layers or at the output layer of the network.
Summary: In this section, we have introduced the most common building blocks used in ML models for short-reach optical communication systems. These fundamental components are still widely used in current approaches. To summarize the complexity of the models discussed earlier, we have provided a table (Table 2) that outlines the complexity analysis for each of the models. In this complexity analysis, we have focused solely on the computation required per batched sample, without considering the choice of hyperparameters like the number of epochs or batch size. This provides a compact overview of the computational demands of each model on a per-sample basis.
@@ 5 Advanced Sequential ML Methods
In Section 4, we introduced traditional sequential models, such as RNN, LSTM, GRU, and CNN. The key question we aim to answer in this section is how to effectively incorporate the unique characteristics of time series data into the modeling process and leverage the temporal convolution model to mitigate channel distortion. Compared to other DL models like transformers and Fourier-based neural networks, convolutional models exhibit better generalization performance. Convolutional models are also more robust to changes in their parameter values when applied to new datasets, unlike the other models which require careful parameter initialization and hyperparameter tuning when used on new data LIU (2022).
This section starts with channel modeling, encompassing four distinct noise models. We derive the characteristics and capabilities required for the algorithm based on these models. Subsequently, we provide a detailed exposition of the architectures and fundamental assumptions underlying three models: Frequency-Calibrated Sampled-Interaction Neural Network (FC-SCINet) CShao (2014)), Light time-series (LightTS) ZhangT (2022) and DLinear Zeng (2022).
5.1 Distortion Model:
The main limiting factor for the equalization task in a short-reach/PON system is ISI as a result of CD, sampling error (jitter), frequency shift (chirp), and Kerr-induced nonlinearity BKaranov (2018). In this section, we will focus on the effects of CD, jitter, and chirp, as these are the dominant distortion mechanisms in short-reach PAM-based systems. The impact of Kerr-nonlinearities is limited in singe-channel PONs due to the relatively short fiber lengths and low optical powers involved.
CD in an optical communication system is caused by different phase velocities with respect to frequency. It fundamentally constitutes a linear transformation, and its mathematical representation involves a differential equation that considers spatial position and time, which can be presented as
@@ 5.1 Distortion Model:
Jitter refers to high-frequency fluctuations in the amplitude of a signal. This high-frequency perturbation can have a significant impact on neural networks that rely on low-frequency signals.
@@
Figure 1: The overview of all sampling modules in temporal convolution networks is modified from CShao (2014); ZhangT (2022), namely interval sampling and continuous sampling in LightTS ZhangT (2022), and interactive sampling in CShao (2014); LIU (2022).
@@ 5.2 Temporal Convolution Neural Network
DL-based equalizers fundamentally capture domain-specific nonlinear disturbances by employing linear transformations and nonlinear activation functions. Within the temporal CNN, the core modules comprise interval, continuous, and interaction sampling modules, alongside convolutional neurons and linear layers, as depicted in Fig. 1. Each of these modules offers practical flexibility for hardware implementation, due to their computational efficiency. Subsequently, we delve into three of the most efficient convolution-based sampling networks.
@@
SCIBlock: The SCIBlock, is a key component of FC-SCINet, because it can iteratively decompose a signal into sub-sequences at various scales, while incorporating nonlinear transformations between adjacent layers. In contrast, the decomp block is restricted to enhancing fixed signal components and is limited to a single scale. From a mathematical perspective, the SCIBlock applies a hierarchical structure by systematically downsampling the input sequence into even-positioned and odd-positioned samples, denoted as $x_{even}$ and $x_{odd}$. Following the convolutional layer, the sub-sequences in adjacent layers are iteratively multiplied together utilizing exponential and multiplication operations, as shown in Eq. (10) and (11).
@@
DLinear: As previously mentioned in FC-SCINet, while the concatenation of the decomp module may not offer optimal equalization, it has exhibited great performance in real-world datasets. Therefore, we will provide a brief introduction to this module: It first decomposes a raw data input into a low $\mathbf{x_{s}}$ and high frequency $\mathbf{x_{f}}$ signal. $\mathbf{x_{s}}$ is extracted by a moving average kernel. It is equivalent to filtering the signal using a $sinc$ function in the frequency domain. These two components are added in a linear combination form, expressed by $W_{s}$, $W_{f}$. The operation above is presented in Eq.(12).
@@
By iterative decomposition with different kernel sizes, DLinear can be extended to a deeper network. The complexity is $O(kn_{s})$, where $k$ is the number of the layer, and $n_{s}$ is the length of the model input. To simplify the complexity, the weight matrix $W$ could be replaced by the convolutional kernel.
@@
Sampling: In contrast to SCIBlock, which samples the raw sequence using odd and even indices, LightTS introduces two generic sampling strategies: Interval Sampling and Continuous Sampling, as shown in Fig. 1. Interval sampling partitions time series data into non-overlapping sub-sequences based on fixed time intervals, as shown in Fig 1. This approach helps identify periodic patterns or regularities within the data while minimizing information loss. On the other hand, continuous sampling divides sequences into corresponding sub-sequences, extracting data points continuously throughout the time series and preserving temporal continuity. This sampling method enables the capture of patterns within the period, ensuring a more comprehensive representation of the underlying dynamics. The subsequent section presents an MLP-based architecture to extract useful features from both the downsampled sub-sequences and continuously sampled sub-sequences.
@@
Information Exchange Block (IEBlock): The IEBlock serves as the central module in LightTS, designed to effectively process the 2D matrix resulting from continuous sampling and interval sampling. This block comprises three essential components: 1) temporal projection, which identifies temporal features following continuous sampling; 2) channel projection, which captures inter-channel information following interval sampling; and 3) the exchange block, which integrates the information from the outputs mentioned above, facilitating information fusion. All of them utilize MLP as the nonlinear behavior learning module. The design of LightTS is notably concise, employing only two sampling modules and an MLP. On certain datasets, it surpasses the performance of FC-SCINet Zeng (2022).
@@
Compared to the models mentioned earlier, the FC-SCINet model requires less structural adaptation and pre-processing when applied in practical PON applications. The FC-SCINet has been successful in recent PON-related work CShao (2014). Different from LSTM, which offers the advantage of ensuring information flow strictly from past to future, temporal CNN goes beyond this by modeling the global dependency between input and output, while also leveraging stacked causal convolution layers. Additionally, the FC-SCINet introduces interaction modeling, enabling the explicit capture of interactions between elements within a sequence, making it a more advanced alternative to LSTM. In addition to these benefits, CNNs and SCINet offer several advantages over LSTM:
@@
  • CNNs can identify patterns regardless of their position within the input sequence. This property makes them well-suited for tasks where the position or timing of relevant features is not fixed, providing greater flexibility compared to LSTM.
@@
  • CNNs excel at capturing local patterns and extracting relevant features from the input sequence. This ability is particularly useful for tasks that require identifying and recognizing specific patterns or motifs within the data.
@@
  • Both CNNs and SCINet architectures typically have fewer parameters compared to LSTM models. This reduced parameter count can make training and inference more efficient, especially when working with limited computational resources or when dealing with large datasets.
@@
Recent Progress: CNNs play a crucial role in current time series prediction research and applications. This is due to their high parameter efficiency, model stability, and strong theoretical foundation (Multiscale Decomposition). The complexity of these convolutional networks mostly depends on the number of layers and the size of the convolutional kernels. Nowadays, more advanced designs like dilated convolution and inception are often combined with other modules to create complex DL models, but they are rarely used on their own. Even so, temporal CNNs still have a distinct advantage in terms of the performance-to-complexity ratio. They are also straightforward to implement in hardware.